# Conformal Quantile Regression for Probabilistic Constitutive Modeling of Anisotropic Soft Materials

### v1

@@ 3.2 Synthetic Data of Arterial Wall Material
For the training, calibration, and testing datasets, we sample 20 realizations from this distribution to generate uniaxial datasets in the axial and circumferential directions, as shown in Figure 8. Although the data are sampled from a homogeneous random process in the model-parameter space, the resulting stress responses exhibit clear heteroscedasticity, as evidenced by the increasing standard deviation with load. This behavior is expected due to the strong nonlinearity of the constitutive model. As one of our goals is to evaluate the extrapolation capability of the proposed scheme, the test data are chosen to extend beyond the maximum stretch observed in the training and calibration datasets.
@@
The synthesized data are available for download at https://....
Figure 8: Training, calibration, and test data. (left) Stress–strain responses for uniaxial tests along different directions. (right) Mean and standard deviation of the data as a function of applied strain.
The hyperparameters used in this example are reported in Section F.2. During training, both axial and circumferential data are used to learn the probabilistic model, and the training curves of the pinball loss computed for each dataset are shown in Figure 9.
@@
Figure 10: Model predictions on the test data (black dots) and uncertainty estimates given by the quantile predictions (dashed orange) and their conformal adjustments (solid purple). Blue points indicate the response corresponding to the mean material parameters of the reference model. The vertical red dashed line denotes the maximum strain used in training.
@@ 3.3 Experiential Data of Porcine Atrioventricular Valve Leaflets
The dataset labeled MVAL1.txt from  [26] is employed. As shown in Fig.˜11, the dataset contains axial and circumferential responses under five different biaxial loading ratios. The first four loading conditions are used for training and calibration, while the fifth is reserved for testing. For each loading scenario except the test case, 80% (randomly selected) of the data are used for training and 20% for calibration. In this figure, one region from the axial stress response and another from the circumferential response at the onset of loading are highlighted. In these regions, the data exhibit considerable scatter, likely due to measurement noise and the highly compliant nature of the material, which makes it sensitive to small load perturbations. Selected regions of the plot are magnified for clarity.
Figure 11: Experimental data of porcine atrioventricular valve leaflets obtained under biaxial loading with varying loading ratios [26]. The test data are shown in red.
As mentioned, data corresponding to four different loading conditions are used to train the quantile model. The hyperparameters and training loss histories are reported in Appendix F.3.
@@
Figure 12: Experimental data of porcine atrioventricular valve leaflets obtained under biaxial loading with varying loading ratios [26]. The test data are shown in red.
@@ 4 Conclusion
While the proposed formulation offers simplicity, ease of training, and fast inference without requiring distributional assumptions, a key limitation is that the resulting conformal prediction intervals primarily capture aleatoric uncertainty associated with data variability. In particular, the approach does not explicitly model epistemic uncertainty arising from model inadequacy or limited data [49]. We will leave addressing this issue for futhre research.
@@ Acknowledgments
Appendix A Proof of Monotonicity
@@
Appendix B Numerical Illustration of Monotone Neural Network
@@
Both models are trained to match sampled gradient data using the mean squared error (MSE) loss. Training is performed for 5000 epochs using the Adam optimizer with a learning rate of $10^{-2}$. To account for variability due to random initialization, each experiment is repeated over $10$ independent runs. The training loss evolution and total training time are reported in Fig.˜13. The monotone parameterization achieves lower training error while requiring approximately half the computational time.
@@
Appendix C Quantile Coverage Guarantee
@@
Appendix D Conformal Marginal Coverage Guarantee
@@
Appendix E loading condition
This appendix summarizes the three canonical incompressible loading modes used to generate and evaluate the constitutive responses: uniaxial tension, equibiaxial and biaxial tension, and pure shear. In all cases we use the deformation gradient $\mathbf{F}$ and enforce incompressibility via $\det(\mathbf{F})=1$ (equivalently $J=1$), with the associated pressure term handled separately in the stress computation.
@@
Uniaxial loading along direction 1 is prescribed by the principal stretches $\lambda_{1}=\lambda,\lambda_{2}=\lambda_{3}=\lambda^{-1/2},$ which yields the deformation gradient $\mathbf{F}_{\mathrm{uni}}(\lambda)=\mathrm{diag}\!\left(\lambda,\lambda^{-1/2},\lambda^{-1/2}\right).$
@@
Biaxial loading in directions 1 and 2 is prescribed by $\lambda_{1},\lambda_{2},\lambda_{3}=(\lambda_{1}\lambda_{2})^{-1},$ leading to $\mathbf{F}_{\mathrm{bi}}(\lambda)=\mathrm{diag}\!\left(\lambda_{1},\lambda_{2},(\lambda_{1}\lambda_{2})^{-1}\right).$
@@
Pure shear is prescribed by opposing stretches in two directions, $\lambda_{1}=\lambda,\lambda_{2}=\lambda^{-1},\lambda_{3}=1,$ and hence $\mathbf{F}_{\mathrm{ps}}(\lambda)=\mathrm{diag}\!\left(\lambda,\lambda^{-1},1\right).$
Appendix F Problem Setting
F.1 Moonery-Rivlin
@@
F.2 Arterial Wall Material
@@
F.3 Porcine Atrioventricular Valve Leaflets
@@
Figure 14: Pinball loss training history for the porcine atrioventricular aalve leaflets data.

### v2

@@ 3.2 Synthetic Data of Arterial Wall Material
For the training, calibration, and testing datasets, we sample 20 realizations from this distribution to generate uniaxial datasets in the axial and circumferential directions, as shown in Figure 9. Although the data are sampled from a homogeneous random process in the model-parameter space, the resulting stress responses exhibit clear heteroscedasticity (i.e., non-constant conditional variance), as evidenced by the increasing standard deviation with load. This behavior is expected due to the strong nonlinearity of the constitutive model. As one of our goals is to evaluate the extrapolation capability of the proposed scheme, the test data are chosen to extend beyond the maximum stretch observed in the training and calibration datasets.
@@
The synthesized data are available for download at https://doi.org/10.21985/n2-bbg3-zq15.
Figure 9: Training, calibration, and test data. (left) Stress–strain responses for uniaxial tests along different directions. (right) Mean and standard deviation of the data as a function of applied strain. Generated data are shown as colored dots, while solid lines denote the corresponding approximated mean and standard deviation.
The hyperparameters used in this example are reported in Section G.2. During training, both axial and circumferential data are used to learn the probabilistic model, and the training curves of the pinball loss computed for each dataset are shown in Figure 10.
@@
More importantly, the probabilistic model exhibits good extrapolation behavior at higher strain levels, closely following the ground-truth response while providing meaningful uncertainty estimates. At very large strains, however, the estimated uncertainty becomes narrower than the variability implied by the data distribution. This behavior reflects operation in an extrapolation regime beyond the maximum strain used for training and calibration. While additional sampling in these regimes would help improve calibration, structural model discrepancy may also contribute to the observed deviations. Moreover, since the present framework primarily captures aleatoric uncertainty, epistemic effects associated with model inadequacy are not explicitly modeled and may become more pronounced in such extrapolation regimes. Since conformal guarantees are marginal and depend on the available calibration data, uncertainty estimates in extrapolation regimes motivate adaptive sampling and potential model refinement to improve coverage.
Figure 11: Model predictions on the test data (light blue dots) and uncertainty estimates given by the quantile predictions (dashed orange lines) and their conformal adjustments (solid black lines). Blue circles indicate the response corresponding to the mean material parameters of the reference model. The vertical red dashed line denotes the maximum strain used in training.
@@ 3.3 Experiential Data of Porcine Atrioventricular Valve Leaflets
The dataset labeled MVAL1.txt from  [30] is employed. As shown in Fig.˜12, the dataset contains axial and circumferential responses under five different biaxial loading ratios. The first four loading conditions are used for training and calibration, while the fifth is reserved for testing. For each loading scenario except the test case, 80% (randomly selected) of the data are used for training and 20% for calibration. In this figure, one region from the axial stress response and another from the circumferential response at the onset of loading are highlighted. In these regions, the data exhibit considerable scatter, likely due to measurement noise and the highly compliant nature of the material, which makes it sensitive to small load perturbations. Selected regions of the plot are magnified for clarity.
Figure 12: Experimental data of porcine atrioventricular valve leaflets under biaxial loading with varying loading ratios, shown in different colors [30]. The test data are highlighted in red. Stress responses along the horizontal direction are denoted by circular markers, while those along the vertical direction are indicated by square markers.
As mentioned, data corresponding to four different loading conditions are used to train the quantile model. The hyperparameters and training loss histories are reported in Appendix G.3.
@@
Figure 13: Results of the proposed method after training on experimental data of porcine atrioventricular valve leaflets under biaxial loading. The data are shown using markers, while the solid lines represent the predicted upper and lower quantiles. Circular markers denote stress responses along the horizontal direction, and square markers denote stress responses along the vertical direction. Red indicates results corresponding to unseen test biaxial loading ratios not included during training and calibration.
@@ 4 Conclusion
While the proposed formulation offers simplicity, ease of training, and fast inference without requiring distributional assumptions, a key limitation is that the resulting conformal prediction intervals primarily capture aleatoric uncertainty associated with data variability. In particular, the approach does not explicitly model epistemic uncertainty arising from model inadequacy or limited data [54]. We will leave addressing this issue for further research.
@@
A structural limitation of the present framework is the assumption of additive separability in the invariant-based energy representation (Eq. 1), which excludes explicit cross-invariant coupling effects. Extending the approach to fully coupled multivariate formulations while preserving admissibility guarantees remains an important direction for future work.
@@ Acknowledgments
Appendix A Supplementary Probabilistic Definitions
This appendix provides a concise summary of key probabilistic concepts used in the manuscript.
A.1 Exchangeability
A sequence $(Z_{1},\dots,Z_{n})$ is said to be exchangeable if
$(Z_{1},\dots,Z_{n})\overset{d}{=}(Z_{\pi(1)},\dots,Z_{\pi(n)})$
for every permutation $\pi$ of $\{1,\dots,n\}$, where $\overset{d}{=}$ denotes equality in distribution.
In the present setting, we apply this notion to the sequence $Z_{i}=(X_{i},Y_{i})$ of input–output pairs. Informally, exchangeability means that the samples are statistically indistinguishable and their ordering carries no information. Independent and identically distributed (i.i.d.) samples form a special case of exchangeable sequences.
A.2 Nonconformity Scores
Conformal prediction calibrates predictive intervals using a data-driven nonconformity score, which measures how much an observation violates a predicted uncertainty set. In the quantile-based setting used in this work, the nonconformity score for a calibration sample $(x_{i},y_{i})$ is defined as
$e_{i}=\max\big(\hat{q}_{\alpha^{\mathrm{lo}}}(x_{i})-y_{i},\;y_{i}-\hat{q}_{\alpha^{\mathrm{hi}}}(x_{i}),\;0\big),$
which quantifies the extent to which the response $y_{i}$ lies outside the predicted quantile interval $\hat{q}_{\alpha^{\mathrm{lo}}}(x_{i}),\hat{q}_{\alpha^{\mathrm{hi}}}(x_{i})]$. If the observation lies within the interval, the score is zero.
A.3 Conformal Adjustment
Let $\{e_{i}\}_{i=1}^{m}$ denote the nonconformity scores computed on the calibration set of size $m$. The conformal adjustment is defined as
$\hat{q}_{\mathrm{adj}}=e_{(k)},\qquad k=\left\lceil(m+1)(1-\alpha)\right\rceil,$
where $e_{(k)}$ denotes the $k$-th smallest calibration score.
A.4 Conformalized Quantile Regression (CQR)
CQR combines quantile regression with conformal calibration to construct predictive intervals with finite-sample marginal coverage guarantees. Given a trained model that predicts conditional quantiles
$\hat{q}_{\alpha^{\mathrm{lo}}}(x),\quad\hat{q}_{\alpha^{\mathrm{hi}}}(x),$
nonconformity scores are computed on a held-out calibration set as
$e_{i}=\max\left\{\hat{q}_{\alpha^{\mathrm{lo}}}(x_{i})-y_{i},\;y_{i}-\hat{q}_{\alpha^{\mathrm{hi}}}(x_{i}),\;0\right\}.$
The conformal adjustment $\hat{q}_{\mathrm{adj}}$ is then defined as the $\left\lceil(m+1)(1-\alpha)\right\rceil$-th smallest nonconformity score among $\{e_{i}\}_{i=1}^{m}$. The final predictive interval is given by
$\left[\hat{q}_{\alpha^{\mathrm{lo}}}(x)-\hat{q}_{\mathrm{adj}},\quad\hat{q}_{\alpha^{\mathrm{hi}}}(x)+\hat{q}_{\mathrm{adj}}\right].$
Appendix B Proof of Monotonicity
@@
Appendix C Numerical Illustration of Monotone Neural Network
@@
Both models are trained to match sampled gradient data using the mean squared error (MSE) loss. Training is performed for 5000 epochs using the Adam optimizer with a learning rate of $10^{-2}$. To account for variability due to random initialization, each experiment is repeated over $10$ independent runs. The training loss evolution and total training time are reported in Fig.˜14. The monotone parameterization achieves lower training error while requiring approximately half the computational time.
@@
Appendix D Quantile Coverage Guarantee
@@
Appendix E Conformal Marginal Coverage Guarantee
@@
Appendix F Loading Condition
This appendix summarizes the three canonical incompressible loading modes used to generate and evaluate the constitutive responses: uniaxial tension, equibiaxial and biaxial tension, and pure shear. In all cases we use the deformation gradient $\bm{F}$ and enforce incompressibility via $\det(\bm{F})=1$ (equivalently $J=1$), with the associated pressure term handled separately in the stress computation.
@@ Appendix F Loading Condition
Uniaxial loading along direction 1 is prescribed by the principal stretches $\lambda_{1}=\lambda,\lambda_{2}=\lambda_{3}=\lambda^{-1/2},$ which yields the deformation gradient $\bm{F}_{\mathrm{uni}}(\lambda)=\mathrm{diag}\!\left(\lambda,\lambda^{-1/2},\lambda^{-1/2}\right).$
@@
Biaxial loading in directions 1 and 2 is prescribed by $\lambda_{1},\lambda_{2},\lambda_{3}=(\lambda_{1}\lambda_{2})^{-1},$ leading to $\bm{F}_{\mathrm{bi}}(\lambda)=\mathrm{diag}\!\left(\lambda_{1},\lambda_{2},(\lambda_{1}\lambda_{2})^{-1}\right).$
@@
Pure shear is prescribed by opposing stretches in two directions, $\lambda_{1}=\lambda,\lambda_{2}=\lambda^{-1},\lambda_{3}=1,$ and hence $\bm{F}_{\mathrm{ps}}(\lambda)=\mathrm{diag}\!\left(\lambda,\lambda^{-1},1\right).$
Appendix G Problem Setting
G.1 Mooney-Rivlin
@@
G.2 Arterial Wall Material
@@
G.3 Porcine Atrioventricular Valve Leaflets
@@
Figure 15: Pinball loss training history for the porcine atrioventricular valve leaflets data.